# thanks: Deceasedthanks: Deceasedthanks: Deceased

### v1

@@ Abstract
A precise characterization of the magnetic properties of LISA Pathfinder free falling test-masses is of special interest for future gravitational wave observatory in space. Magnetic forces have an important impact on the instrument sensitivity in the low frequency regime below the millihertz. In this paper we report on the magnetic injection experiments performed throughout LISA Pathfinder operations. We show how these experiments allowed a high precision estimate of the instrument magnetic parameters. The remanent magnetic moment was found to have a modulus of $(0.245\pm 0.081)\,\rm{nAm}^{2}$, the x-component of the background magnetic field within the test masses position was measured to be $(414\pm 74)$ nT and its gradient had a value of $(-7.4\pm 2.1)\,\mu$T/m. Finally, we also measured the test mass magnetic susceptibility to be (3.35±0.15)×(-3.35$\pm$ 0.15$\times$( - 3.35 ± 0.15 ) ×10-5 in the low frequency regime. All results are in agreement with on-ground estimates.
@@ I Introduction
LISA Pathfinder (LPF) [1, 2] was an ESA mission designed as a technology demonstrator for the future gravitational wave observatory in space, LISA [3]. The main goal of the mission was to demonstrate key technologies required to detect gravitational waves in space. In order to do so, the instrument on-board had to achieve a relative acceleration noise between its two test masses (TMs) in nominal geodesic motion down to $\rm 3\times 10^{-14}\,ms^{-2}Hz^{-1/2}$ at 1 mHz. The relevance of the measurements being its precision in the very low frequency band, a region not achievable with on ground gravitational wave detectors.
LPF launched on December 3rd, 2015 and started its scientific operations on the 1st of March of 2016 after reaching the Lagrange point L1 of the Earth-Sun system. The mission was divided into two different experiments on-board, the European LISA Technology Package (LTP) and the American Disturbance Reduction System (DRS). After seventeen months of scientific operations, the mission successfully demonstrated its main scientific goal, even achieving values of the residual acceleration power spectral density below the LISA requirements along the entire measuring frequency band [20, 4].
Achieving such a demanding level of geodesic free fall was as important as understanding all the different contributions that build the noise model of the instrument. With the goal to design a suitable model for future space-born missions, several experiments were planned during the LPF operations with the scope to split up the noise measurements into its main contributions. With that objective, LISA Pathfinder carried the Data and Diagnostics Subsystem (DDS), which included a temperature measurement subsystem [5, 6], a magnetic diagnostic subsystem [7, 8] and a radiation monitor [9, 10, 11, 12].
In this work we will focus on the results of the magnetic diagnostics and, specifically, on the experiments run to characterise the magnetic parameters of the test masses on-board LPF. Getting to know such values more precisely is crucial for the future space-borne gravitational wave observatories, since any magnetic perturbation can have a potential impact on the instrument performance through magnetic parasitic forces.
This work is organised as follows. In section II we describe the magnetic diagnostic system on-board designed to study and disentangle the nature of the magnetic forces, introduced in section III, that can perturb the test mass motion. In section IV we describe the in-flight magnetic experiments performed to extract the TMs magnetic parameters and we present our conclusions in section V.
@@ II.1 The magnetic diagnostics subsystem
The magnetic diagnostics subsystem on-board LISA Pathfinder was in charge of monitoring the magnetic environment and, as well, to induce controlled magnetic fields to perturb the test masses motion in order to properly characterise the contribution of magnetic forces to the total instrument noise budget. To achieve these goals, the subsystem was composed by four tri-axial magnetometers and two induction coils.
The coils —see Fig. 1— were able to inject a controlled magnetic field within the TMs position as well as in the magnetometers closest to them. Both circular induction coils, with an average radius of 56.5 mm, were located 85.5 mm away from the test masses and they were attached to the external wall of each vacuum enclosure. The wire winding used to build up the coils were made of a Titanium alloy ($Ti_{6}Al_{4}V$) and loop around the structure for a total of 2400 turns. Both coils centres were aligned with the axis, x, joining both TMs centres so that the induced magnetic field had axial symmetry. The four magnetometers were aligned by pairs in each one of the spacecraft axes in order to be able to measure gradients within the spacecraft in both the y and x directions. The entire system was set-up at a same height in the z axis, meaning that no magnetic field gradients along such direction could be measured. Magnetometers were continuously measuring with high precision the evolution of the on-board magnetic field. For each of the measuring axes, each fluxgate magnetometer had a sensing coil surrounding a second inner drive coil around a high permeability magnetic core material. This meaning that all four magnetometers contained active magnetic sensors that had to be located far enough from the test masses for them not to contribute as a source of magnetic parasitic forces.
@@ II.2 Magnetic environment on-board
The background magnetic field measured on-board was completely dominated by the contribution from the electronics of the spacecraft units. Among them, the thruster systems were a major contributor, both the cold gas high pressure latch valves (the ones used by ESA) and the colloidal thrusters (the ones operated by NASA). Cold gas thrusters or, more precisely, some permanent magnets in the cold gas thruster subsystem, contributed with roughly the 80% of the measured magnetic field. Although a strong contribution, this one remained constant throughout the mission – partially thanks to the high thermal stability reached on-board [6] – which is key for a mission as LISA with strong requirements on any potential source of fluctuations. This is not the case for the colloidal thrusters, where a persistent slow drift of around 150  nT in the span of 100 days was observed [8]. The main contribution of the magnetic-induced force noise is below the millihertz [13]. In this frequency regime, magnetic field fluctuations are dominated by the interplanetary magnetic field contribution, which can show an important non-stationary component associated with changes in the interplanetary plasma. For instance, variations in the range of $\rm 300-500\,km\,s^{-1}$ in the solar wind velocity were related to variations in the amplitude spectral density in the range $\rm 20-50\,\mu Hz$ of around $\rm 170-750\,nT\ Hz^{-1/2}$ [8].
In what refers to our analysis in the following, we can safely assume that in all the in-flight experiments where we induced magnetic fields with the coils, the background magnetic field (either generated by the spacecraft or due to the interplanetary contribution) can be safely neglected as it was at least one order of magnitude smaller than the ones induced by the coils. The same is true for the gradients of the magnetic fields.
@@ III.1 A magnetic dipole in a surrounding magnetic field
In a first approximation, the free-falling test masses inside LISA Pathfinder can be considered as a magnetic dipole with total magnetic moment density $\mathbf{m}$ inside of a surrounding magnetic field $\mathbf{B}$. The dipole would therefore feel an associated force and torque given by
@@
where $\mathbf{r}$ denotes the distance to the TM with respect to the coil and we also use the convention $\left\langle\ldots\ \right\rangle\equiv\dfrac{1}{\rm{V}}\int_{\rm{V}}\,(\ldots)\,d^{3}x$ to denote TM volume, V, average of the enclosed quantity. In our notation, parameters in bold refer to vectors. The total magnetic moment density $\mathbf{m}$ is the sum of two components: the remanent magnetic moment density $\mathbf{m_{r}}$ ($\mathbf{m}$ from now on) which depends on the material and manufacturing process and the induced magnetic moment density, $\mathbf{m_{i}}$. Given that the dominant material in the test mass composition (73% Au, 27% Pt) is gold, we assume the magnetic moment to be proportional to the applied magnetic field, i.e.
@@
where $\chi$ is the magnetic susceptibility. Notice that, although we expect a typical negative susceptibility from a diamagnetic material, we leave the sign undetermined in the following derivation. Also, we have implicitly assumed here an isotropic test mass which allows a scalar susceptibility in the previous equation. Next, we assume the test mass magnetized by a low alternating magnetic field
@@
where $\mathbf{B}^{AC}$ is the amplitude of the oscillating magnetic field and $\omega$ the frequency of the applied magnetic field, we will obtain a magnetization that varies with time accordingly, $\mathbf{m_{i}}(t)$. In diamagnetic, paramagnetic and many ferromagnetic materials, the magnetization also varies sinusoidally and in phase with the applied magnetic field with a constant ratio given by the magnetic susceptibility. However, some ferromagnetic materials show a delayed response that is not in phase with the applied field. This phenomena is typically described by considering the in-phase ($\chi_{r}$) and out-of-phase ($\chi_{i}$) components of the magnetic susceptibility. For the case of LPF experiments the most relevant physical mechanism involved in the latter one are Eddy Currents since this contribution becomes increasingly important with the increasing conductivity of the material. However, for most of the experiments at the low frequency regime this component is expected to be orders of magnitude smaller thus, its contribution can be neglected.
Considering both contributions of the magnetization (the remanent and the induced magnetic moments) one can further develop Eqs. (1) into
$\mathbf{F}$ $=\left\langle\left(\mathbf{m}\cdot\mathbf{\nabla}\right)\mathbf{B}+\frac{\chi}{\mu_{0}}\,\left[\left(\mathbf{B}\cdot\mathbf{\nabla}\right)\mathbf{B}\right]\right\rangle\rm{V},$ (4a)
$\mathbf{N}$ $=\left\langle\mathbf{m}\times\mathbf{B}+\mathbf{r}\times\left[(\mathbf{m}\cdot\nabla)\,\mathbf{B}+\frac{\chi}{\mu_{0}}(\mathbf{B}\cdot\mathbf{\nabla})\,\mathbf{B}\right]\right\rangle\rm{V}.$ (4b)
In order to describe our experiments in the following sections, we still need to include some more degree of detail in the equations. First, we will consider the magnetic field as composed by an applied, oscillating magnetic field $\mathbf{B}^{AC}$, and a stable magnetic field $\mathbf{B}_{0}$, divided into an applied time independent DC magnetic field $\mathbf{B}^{DC}$ and some environmental background $\mathbf{B}_{back.}$
@@
$\begin{split}\mathbf{F_{DC}}=\Biggl{[}&\left\langle\left(\mathbf{M}\cdot\mathbf{\nabla}\right)\mathbf{B}_{0}\right\rangle\\ &+\frac{\chi\rm{V}}{\mu_{0}}\left(\left\langle\left(\mathbf{B_{0}}\cdot\mathbf{\nabla}\right)\mathbf{B_{0}}\right\rangle+\frac{1}{2}\left\langle(\mathbf{B}^{AC}\cdot\mathbf{\nabla})\mathbf{B}^{AC}\right\rangle\right)\Biggr{]},\\ \end{split}$ (7a)
$\begin{split}\mathbf{F_{1\omega}}=\Biggl{[}&\left\langle\left(\mathbf{M}\cdot\mathbf{\nabla}\right)\mathbf{B}^{AC}\right\rangle\\ &+\frac{\chi\rm{V}}{\mu_{0}}\biggl{(}\left\langle\left(\mathbf{B}_{0}\cdot\mathbf{\nabla}\right)\mathbf{B}^{AC}\right\rangle+\left\langle\left(\mathbf{B}^{AC}\cdot\mathbf{\nabla}\right)\mathbf{B}_{0}\right\rangle\biggr{)}\Biggr{]}\\ &\times\sin(\omega\,t),\\ \end{split}$ (7b)
@@
where $\mathbf{M}=\mathbf{m}\rm{V}$ is the remanent magnetic moment and we have assumed homogeneity and stationarity of the test mass properties. Considering that the relative acceleration measurements in LISA Pathfinder are in the x direction, the only component of the force that will be needed is its x component. Analogously, if we manipulate the torque equations a similar result with the three terms before mentioned should appear but that won’t be the case as we will see in the next section.
@@ III.2 Estimate of test mass magnetic parameters
The coupling between an induced magnetic field and its gradient with the magnetic susceptibility of the test mass is responsible of the appearance of a force component at twice the injected modulation frequency in Eq. (7c). Our analysis can take advantage of this by extracting the signal at $2\omega$ from the measured test mass force, i.e.
$\chi_{2\omega}=-\frac{2\,\mu_{0}}{V}\frac{\hat{F}_{2\omega,x}}{\left\langle B_{x}^{AC}\right\rangle\cdot\left\langle\nabla_{x}B_{x}^{AC}\right\rangle},$ (8)
This equation provides a direct estimate of the test mass susceptibility decoupled from any other of the magnetic parameters. The notation in Eq. (8) shows explicitly that the estimate of the susceptibility is obtained at twice the injected frequency by demodulating the encoded information in the $\Delta g$ and comparing it with the predicted TM average magnetic field and gradient.
@@
The component of the force at the injection frequency, $F_{1\omega,x}$, mixes all the parameters we might be interested in. Taking into advantage that all terms in Eq. (7b) depend on $\left\langle B_{x}^{AC}\right\rangle$ or $\left\langle\nabla_{x}B_{x}^{AC}\right\rangle$, we can rewrite the expression as follows
$\hat{F}_{1\omega,x}=\left[M_{x}+\frac{\chi V}{\mu_{0}}(B_{0,x}+\kappa\nabla_{x}B_{0,x})\right]\left\langle\nabla_{x}B_{x}^{AC}\right\rangle.$ (9)
@@
$\begin{split}M_{eff,x}=M_{x}+\frac{\chi V}{\mu_{0}}\Biggl{[}&\left(B_{back.,x}+\left\langle B_{x}^{DC}\right\rangle\right)\\ &+\kappa\nabla_{x}\left(B_{back.,x}+\left\langle B_{x}^{DC}\right\rangle\right)\Biggr{]},\end{split}$ (10)
where we have expanded $B_{0,x}$ as explained in Eq. (5). $B_{back.,x}$ can be considered negligible in front of the injected magnetic field $\left\langle B_{x}^{DC}\right\rangle$ as its value is expected to be an order of magnitude smaller than the injected fields through the coils. Thus, we end up having
$M_{\text{eff},x}\simeq M_{x}+\frac{2\,\chi V}{\mu_{0}}\left\langle B_{x}^{DC}\right\rangle.$ (11)
If the only variable in Eq. (11) is $\left\langle B_{x}^{DC}\right\rangle$, we will obtain a straight line with an offset that corresponds to the remanent magnetic moment $M_{x}$ and a slope that is proportional to the magnetic susceptibility $\chi$ at $1\omega$. Furthermore, when we induce a magnetic field in the TM position, using the coils, apart from direct forces in the x direction, we are also generating torques, as described in Eq. (4b). Due to the symmetry of the system the term involving the cross product with $\mathbf{r}$ will integrate to zero accross the TM volume due to the alignment between the coil axis and the TMs center resulting in only two components, see Appendix B for the explanation, where only the $1\omega$ term will be of interest leading to the following equations
$\hat{N}_{\phi,1\omega}=-M_{y}\left\langle B_{x}^{AC}\right\rangle\ \ ;\ \ \hat{N}_{\eta,1\omega}=M_{z}\left\langle B_{x}^{AC}\right\rangle.$ (12)
We can conclude that the $1\omega$ oscillation of the torque in $\phi$ is directly related to the remanent magnetic moment along y, $M_{y}$, while the $1\omega$ oscillation of the torque in $\eta$ is directly related to the remanent magnetic moment along z, $M_{z}$. Therefore, by demodulating the torque at $1\omega$ for $\eta$ and $\phi$ and considering the values of the injected magnetic field $\left\langle B_{x}^{AC}\right\rangle$, we will be able to determine $M_{y}$ and $M_{z}$.
@@
Similarly to the $1\omega$ term, in Eq. (7a), the expression of the DC force component of the signal involves again all the unknown parameters. If we group all the terms of Eq. (7a) as a function of $\left\langle B_{x}^{DC}\right\rangle$, we can rewrite it as:
$\begin{split}&\hat{F}_{DC,x}\simeq\left(\frac{\chi V}{\mu_{0}\kappa}\right)\left\langle B_{x}^{DC}\right\rangle^{2}+\\ \end{split}$
$\begin{split}+\left[\frac{M_{x}}{\kappa}+\frac{\chi V}{\mu_{0}}\left(\nabla_{x}B_{back.,x}+\frac{B_{back.,x}}{\kappa}\right)\right]\left\langle B_{x}^{DC}\right\rangle+\\ \end{split}$
@@
where $M_{+}=M_{x}+M_{y}+M_{z}$. We have made the assumption that the background magnetic field is the same in all directions, $B_{back.,x}\simeq B_{back.,y}\simeq B_{back.,z}$, because we don’t have any information a priori on the magnitude of such variable. We also had to consider that the gradient of $B_{back.,x}$ was the same along all three axes: $\nabla_{x}B_{back.,x}\simeq\nabla_{y}B_{back.,x}\simeq\nabla_{z}B_{back.,x}$, which is a worst case scenario since all components contributing to the background gradient would add up when in reality they could cancel each other. If the only variable is $\left\langle B_{x}^{DC}\right\rangle$, we can observe $\hat{F}_{DC,x}$ follows a quadratic equation.

### v2

@@ Abstract
A precise characterization of the magnetic properties of LISA Pathfinder free falling test-masses is of special interest for future gravitational wave observatory in space. Magnetic forces have an important impact on the instrument sensitivity in the low frequency regime below the millihertz. In this paper we report on the magnetic injection experiments performed throughout LISA Pathfinder operations. We show how these experiments allowed a high precision estimate of the instrument magnetic parameters. The remanent magnetic moment was found to have a modulus of $(0.245\pm 0.081)\,\rm{nAm}^{2}$, the x-component of the background magnetic field within the test masses position was measured to be $(414\pm 74)$ nT and its gradient had a value of $\rm(-7.4\pm 2.1)\,\mu$T/m. Finally, we also measured the test mass magnetic susceptibility at 5 mHz to be (3.3723±0.0069)×(-3.3723$\pm$ 0.0069$\times$( - 3.3723 ± 0.0069 ) ×10-5. All results are in agreement with on-ground estimates.
@@ I Introduction
LISA Pathfinder (LPF) [1, 2] was an ESA mission designed as a technology demonstrator for the future gravitational wave observatory in space, LISA [3]. The main goal of the mission was to demonstrate key technologies required to detect gravitational waves in space. In order to do so, the instrument on-board had to demonstrate a relative acceleration noise between its two test masses (TMs) in nominal geodesic motion at a level of $\rm 3\times 10^{-14}\,ms^{-2}Hz^{-1/2}$ at 1 mHz, a level of precision impossible to achieve with on ground gravitational wave detectors.
LPF launched on December $\rm 3^{rd}$, 2015 and started its scientific operations on the March $\rm 1^{st}$, 2016 after reaching the Lagrange point L1 of the Earth-Sun system. The mission was divided into two different experiments on-board, the European Space Agency LISA Technology Package (LTP) and the NASA Disturbance Reduction System (DRS). After seventeen months of scientific operations, the mission successfully demonstrated its main scientific goal, surpassing its requirements and achieving a level of acceleration noise below the LISA requirements in its entire measurement frequency band [4, 5].
As important as achieving this demanding level of geodesic free fall was the development of an understanding all the different contributions that build the noise model of the instrument. Several experiments were planned during the LPF operations in order to isolate and evaluate the most important contributions to the acceleration noise budget. With that objective, LISA Pathfinder carried the Data and Diagnostics Subsystem (DDS), which included a temperature measurement subsystem [6, 7], a magnetic diagnostic subsystem [8, 9] and a radiation monitor [10, 11, 12, 13].
In this work we will focus on the results of the magnetic diagnostics and, specifically, on the experiments run to characterize the magnetic parameters of the test masses on-board LPF. Precise knowledge of such values is crucial for the future space-borne gravitational wave observatories, since any magnetic perturbation can have a potential impact on the instrument performance through magnetic parasitic forces.
This work is organized as follows. In section II we describe the magnetic diagnostic system on-board designed to study and disentangle the nature of the magnetic forces, introduced in section III, that can perturb the test mass motion. In section IV we describe the in-flight magnetic experiments performed to extract the TMs magnetic parameters and we present our conclusions in section V.
@@ II.1 The magnetic diagnostics subsystem
The magnetic diagnostics subsystem on-board LISA Pathfinder was responsible for monitoring the magnetic environment and creating controlled magnetic fields to perturb the test mass motion in order to properly characterize the contribution of magnetic forces to the total instrument noise budget. To achieve these goals, the subsystem was composed by four triaxial magnetometers and two induction coils.
The coils—see Fig. 1—were able to produce a controlled magnetic field at the TM locations as well as in the magnetometers closest to them. Both circular induction coils, with an average radius of 56.5 mm, were located 85.5 mm away from the test masses and they were attached to the external wall of each vacuum enclosure. The wire winding used to build up the coils were made of a Titanium alloy (Ti6Al4V) and loop around the structure for a total of 2400 turns. The centers of both coils were aligned with the axis, $x$, joining both TMs centers so that the induced magnetic field had axial symmetry. The four magnetometers were aligned by pairs in the $x-y$ plane of the TMs in order to be able to measure gradients within the spacecraft in both the $x$ and $y$ directions. Magnetic field gradients along the $z$ direction could not be measured. The magnetometers continuously measured the evolution of the on-board magnetic field with a precision of 10 nT Hz-1/2. Each fluxgate magnetometer measurement axis consisted of a sensing coil surrounding a second inner drive coil around a high permeability magnetic core material. This meant that all four magnetometers contained active magnetic sensors that had to be located far enough from the test masses for them not to contribute as a source of magnetic parasitic forces.
@@ II.2 Magnetic environment on-board
The background magnetic field measured on-board was completely dominated by the contribution from the electronics of the spacecraft units. Among them, the thruster systems were a major contributor, both the cold gas high pressure latch valves (the ones used by ESA) and the colloidal thrusters (the ones operated by NASA). Cold gas thrusters or, more precisely, some permanent magnets in the cold gas thruster subsystem, contributed with roughly the 80% of the measured magnetic field. Although a strong contribution, this remained constant throughout the mission—partially thanks to the high thermal stability reached on-board [7]. This was not the case for the colloidal thrusters, where a persistent slow drift of around 150 nT in the span of 100 days was observed [9]. The main contribution of the magnetic-induced force noise is below the millihertz [14]. In this frequency regime, magnetic field fluctuations are dominated by the interplanetary magnetic field contribution, which can show an important non-stationary component associated with changes in the interplanetary plasma. For instance, variations in the range of $\rm 300-500\,km\,s^{-1}$ in the solar wind velocity were found to be correlated to variations in the magnetic fields amplitude spectral density in the range $\rm 20-50\,\mu Hz$ of around $\rm 170-750\,nT\ Hz^{-1/2}$ [9].
In what refers to our analysis in the following, we assume that in all the in-flight experiments where we induced magnetic fields with the coils, the background magnetic field (either generated by the spacecraft or due to the interplanetary contribution) can be safely neglected as it was at least one order of magnitude smaller than the ones induced by the coils. The same is true for the gradients of the magnetic fields.
II.3 Forces on-board LPF
We evaluate the induced force in the test mass through the $\Delta g$ variable—the principal scientific output of the mission—nominally defined as the differential acceleration between the two TMs in their nominal position [4]. Since the main objective of the $\Delta g$ measurement is the evaluation of the free fall of the test masses, those forces arising due to the spacecraft dynamics control loop or other forces caused by spacecraft non-inertial reference frame are subtracted in the definition of this parameter [15]. We will assume that during magnetic injections the dominant forces that TMs will feel will be purely of magnetic origin. $\Delta g$, by construction, is the difference in acceleration between TMs along the axis joining them, the same as the spacecraft $x$ axis by definition and aligned with the $x$ axis from Fig. 1,
$\Delta g=\dfrac{F_{2,x}}{M_{\rm{TM1}}}-\dfrac{F_{1,x}}{M_{\rm{TM2}}},$ (1)
where $M_{\rm{TM}}$ is the mass of each TM and $F$ is the total force each one feels. When calculating the force measured by each TM during magnetic injections, we assume that the TM furthest from the injecting coil will feel a negligible force when compared to the one nearest to the coil. So, when injecting magnetic signals with the coils, we will estimate the magnetic force that the nearest TM feels as $F_{x}\equiv\rm{M_{TM}}\Delta g$.
The same is true for the torque such that $N_{\eta,\phi}\equiv I\Delta g_{\eta,\phi}$, where $I$ is the moment of inertia of a cube and $\Delta g_{\eta,\phi}$ are the angular accelerations measured by the interferometric system. Only rotations along the $y$ and $z$ axes, $\eta$ and $\phi$ respectively in Fig 1, are obtainable since the instrument is not sensitive to rotations around the $x$ axis as it is aligned with the laser beam.
@@ III.1 A magnetic dipole in a surrounding magnetic field
In a first approximation, the free-falling test masses inside LISA Pathfinder can be considered as a magnetic dipole with total magnetic moment density $\mathbf{m}$ inside of a surrounding magnetic field $\mathbf{B}$. Parameters in bold refer to vectors. The dipole would therefore feel an associated force and torque given by
@@
where $\mathbf{r}$ denotes the distance to the TM with respect to the coil. We use the convention $\left\langle\ldots\ \right\rangle\equiv\dfrac{1}{\rm{V}}\int_{\rm{V}}\,(\ldots)\,d^{3}x$ to denote the average of the enclosed quantity over the TM volume, V. The total magnetic moment density $\mathbf{m}$ is the sum of two components: the remanent magnetic moment density $\mathbf{m_{r}}$ which depends on the material and manufacturing process and the induced magnetic moment density, $\mathbf{m_{i}}$. The induced magnetic dipole density of any material is proportional to the applied magnetic field, i.e.
@@
where $\chi$ is the magnetic susceptibility and $\mu_{0}$ is the vacuum permeability. The test mass composition is 73% gold, diamagnetic, and 27% platinum, paramagnetic. Given that the dominant material is gold we should expect the TM to behave like any diamagnetic material, opposing the external magnetic field thus, having a negative magnetic susceptibility. Despite this, we leave the sign undetermined in the following derivation. Also, we have implicitly assumed here an isotropic test mass which allows the use of a scalar susceptibility in the previous equation. Next, we assume the test mass magnetized by a slowly oscillating field
@@
where $\mathbf{B}^{AC}$ is the amplitude of the oscillating magnetic field and $\omega$ its angular frequency, we will obtain a magnetization that varies with time accordingly, $\mathbf{m_{i}}(t)$. In diamagnetic, paramagnetic and many ferromagnetic materials, the magnetization also varies sinusoidally and in phase with the applied magnetic field with a constant ratio given by the magnetic susceptibility. However, some ferromagnetic materials show a delayed response that is not in phase with the applied field. This phenomena is typically described by considering the in-phase or real, $\chi_{r}$, and the out-of-phase or imaginary, $\chi_{i}$, components of the magnetic susceptibility such that $\chi=\chi_{r}+i\chi_{i}$, where $i=\sqrt{-1}$. For the case of LPF experiments the most relevant physical mechanism involved in the imaginary susceptibility are eddy currents since this contribution becomes increasingly important with increasing conductivity of the material. However, for most of the experiments in the low frequency regime this component is expected to be orders of magnitude smaller than the real part thus, its contribution can be neglected. In the following, therefore, we refer to the magnetic susceptibility $\chi$ as equivalent to its real component $\chi_{r}$, unless explicitly stated otherwise.
Considering both contributions of the magnetization (the remanent and the induced magnetic moments) we expand Eqs. (2) into
$\mathbf{F}$ $=\left\langle\left(\mathbf{m_{r}}\cdot\mathbf{\nabla}\right)\mathbf{B}+\frac{\chi}{\mu_{0}}\,\left[\left(\mathbf{B}\cdot\mathbf{\nabla}\right)\mathbf{B}\right]\right\rangle\rm{V},$ (5a)
$\mathbf{N}$ $=\left\langle\mathbf{m_{r}}\times\mathbf{B}+\mathbf{r}\times\left[(\mathbf{m_{r}}\cdot\nabla)\,\mathbf{B}+\frac{\chi}{\mu_{0}}(\mathbf{B}\cdot\mathbf{\nabla})\,\mathbf{B}\right]\right\rangle\rm{V}.$ (5b)
In order to describe our experiments in the following sections, we need to develop the equations further. First, we will consider the magnetic field as composed by an applied, oscillating magnetic field $\mathbf{B}^{AC}$, and a stable magnetic field $\mathbf{B}_{0}$, divided into an applied time independent DC magnetic field $\mathbf{B}^{DC}$ and some environmental background $\mathbf{B}_{back.}$
@@
$\begin{split}\mathbf{F_{DC}}=\Biggl{[}&\left\langle\left(\mathbf{M_{r}}\cdot\mathbf{\nabla}\right)\mathbf{B}_{0}\right\rangle\\ &+\frac{\chi\rm{V}}{\mu_{0}}\left(\left\langle\left(\mathbf{B_{0}}\cdot\mathbf{\nabla}\right)\mathbf{B_{0}}\right\rangle+\frac{1}{2}\left\langle(\mathbf{B}^{AC}\cdot\mathbf{\nabla})\mathbf{B}^{AC}\right\rangle\right)\Biggr{]},\\ \end{split}$ (8a)
$\begin{split}\mathbf{F_{1\omega}}=\Biggl{[}&\left\langle\left(\mathbf{M_{r}}\cdot\mathbf{\nabla}\right)\mathbf{B}^{AC}\right\rangle\\ &+\frac{\chi\rm{V}}{\mu_{0}}\biggl{(}\left\langle\left(\mathbf{B}_{0}\cdot\mathbf{\nabla}\right)\mathbf{B}^{AC}\right\rangle+\left\langle\left(\mathbf{B}^{AC}\cdot\mathbf{\nabla}\right)\mathbf{B}_{0}\right\rangle\biggr{)}\Biggr{]}\\ &\times\sin(\omega\,t),\\ \end{split}$ (8b)
@@
where $\mathbf{M_{r}}=\mathbf{m_{r}}\rm{V}$ is the remanent magnetic moment and we have assumed homogeneity and stationarity of the test mass properties. Considering that the relative acceleration measurements in LISA Pathfinder are in the $x$ direction, the only component of the force from Eq. 7 that will be needed is its $x$ component. Analogously, if we manipulate the torque equations a similar result with the three terms before mentioned should appear.
@@ III.2 Estimate of test mass magnetic parameters
The coupling between an induced magnetic field and its gradient with the magnetic susceptibility of the test mass is responsible for the appearance of a force component at twice the injected modulation frequency in Eq. (8c). Our analysis can take advantage of this by extracting the signal at $2\omega$ from the measured test mass force, i.e.
$\chi_{2\omega}=-\frac{2\,\mu_{0}}{V}\frac{\hat{F}_{2\omega,x}}{\left\langle B_{x}^{AC}\right\rangle\cdot\left\langle\nabla_{x}B_{x}^{AC}\right\rangle}.$ (9)
This equation provides a direct estimate of the test mass susceptibility decoupled from any other of the magnetic parameters. The notation in Eq. (9) shows explicitly that the estimate of the susceptibility is obtained at twice the injected frequency by demodulating the encoded information in $\Delta g$ and comparing it with the predicted TM average magnetic field and gradient.
@@
The component of the force at the injection frequency, $F_{1\omega,x}$, depends on all the magnetic parameters of the TM. Taking advantage of the fact that all terms in Eq. (8b) depend on $\left\langle B_{x}^{AC}\right\rangle$ or $\left\langle\nabla_{x}B_{x}^{AC}\right\rangle$, we can rewrite the expression as follows
$\hat{F}_{1\omega,x}=\left[M_{r,x}+\frac{\chi V}{\mu_{0}}(B_{0,x}+\kappa\nabla_{x}B_{0,x})\right]\left\langle\nabla_{x}B_{x}^{AC}\right\rangle.$ (10)
@@
$\begin{split}M_{\text{eff}f,x}=M_{r,x}+\frac{\chi V}{\mu_{0}}\Biggl{[}&\left(B_{back.,x}+\left\langle B_{x}^{DC}\right\rangle\right)\\ &+\kappa\nabla_{x}\left(B_{back.,x}+\left\langle B_{x}^{DC}\right\rangle\right)\Biggr{]},\end{split}$ (11)
where we have expanded $B_{0,x}$ as explained in Eq. (6). $B_{back.,x}$ can be considered negligible compared to the injected magnetic field $\left\langle B_{x}^{DC}\right\rangle$ as its value is expected to be an order of magnitude smaller than the injected fields through the coils. Thus, we have
$M_{\text{eff},x}\simeq M_{r,x}+\frac{2\,\chi V}{\mu_{0}}\left\langle B_{x}^{DC}\right\rangle.$ (12)
If the only variable in Eq. (12) is $\left\langle B_{x}^{DC}\right\rangle$, then plotting Eq. (10), we will obtain a straight line with an offset that corresponds to the remanent magnetic moment $M_{r,x}$ and a slope that is proportional to the magnetic susceptibility $\chi$ at $1\omega$. Furthermore, when we induce a magnetic field in the TM position, using the coils, apart from direct forces in the $x$ direction, we are also generating torques, as described in Eq. (5b). Due to the symmetry of the system the term involving the cross product with $\mathbf{r}$ will integrate to zero accross the TM volume due to the alignment between the coil axis and the TMs center resulting in only two components, see Appendix B for details, where only the $1\omega$ term will be of interest leading to the following equations
$\hat{N}_{\phi,1\omega}=-M_{r,y}\left\langle B_{x}^{AC}\right\rangle\ \ ;\ \ \hat{N}_{\eta,1\omega}=M_{r,z}\left\langle B_{x}^{AC}\right\rangle.$ (13)
We can conclude that the $1\omega$ oscillation of the torque in $\phi$ is directly related to the remanent magnetic moment along $y$, $M_{r,y}$, while the $1\omega$ oscillation of the torque in $\eta$ is directly related to the remanent magnetic moment along $z$, $M_{r,z}$. Therefore, by demodulating the torque at $1\omega$ for $\eta$ and $\phi$ and considering the values of the injected magnetic field $\left\langle B_{x}^{AC}\right\rangle$, we will be able to determine $M_{r,y}$ and $M_{r,z}$.
@@
Similarly to the $1\omega$ term, in Eq. (8a), the expression of the DC force component of the signal involves again all the unknown parameters. If we group all the terms of Eq. (8a) as a function of $\left\langle B_{x}^{DC}\right\rangle$, we can rewrite it as
$\begin{split}&\hat{F}_{DC,x}\simeq\left(\frac{\chi V}{\mu_{0}\kappa}\right)\left\langle B_{x}^{DC}\right\rangle^{2}\\ \end{split}$
$\begin{split}+\left[\frac{M_{r,x}}{\kappa}+\frac{\chi V}{\mu_{0}}\left(\nabla_{x}B_{back.,x}+\frac{B_{back.,x}}{\kappa}\right)\right]\left\langle B_{x}^{DC}\right\rangle\\ \end{split}$
@@
where $M_{+}=M_{r,x}+M_{r,y}+M_{r,z}$. We have made the assumption that the background magnetic field is the same in all directions, $B_{back.,x}\simeq B_{back.,y}\simeq B_{back.,z}$, because we don’t have any a priori information about its value. We also assume that the three components of $\nabla B_{back.,x}$ are equal: $\nabla_{x}B_{back.,x}\simeq\nabla_{y}B_{back.,x}\simeq\nabla_{z}B_{back.,x}$, which is a worst case scenario since all components contributing to the background gradient would add up when in reality they could partially cancel each other. If the only variable is $\left\langle B_{x}^{DC}\right\rangle$, we can observe that $\hat{F}_{DC,x}$ has a quadratic form.